Session: Nov 1997 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (orfei at Tue Nov 11 07:02:36 MET 1997) Success
No problems encountered

Westerbork (tony foley at Tue Nov 11 11:13:43 MET 1997) Minor failures
No phasecal Unexplained drop in TP IF1 (=odd VCs) from about 1500UT, after which the label was stable at a lower level

Onsala (A. Polatidis at Tue Nov 11 12:04:05 MET 1997) Success
No known problems. We only record LCP so expect signal on ODD VCs only.

Torun (K.Borkowski at Tue Nov 11 12:48:31 MET 1997) Minor failures
Loss of the telescope control (thus probably off source part of the time) between 3:40 and 3:50 UT on Nov. 11. Otherwise all seems O.K.

Jodrell (Alastair Gunn at Wed Nov 12 12:01:06 MET 1997) Minor failures
Late on-source at 17:00. Telescope azimuth limit reached and data lost for several scans during unwrap at about 01:00 UT. Track 21 dead. No other known problems.

Effelsberg (dgraham at Thu Nov 13 11:09:36 MET 1997) Success

Noto (C.Trigilio at Fri Nov 14 09:37:41 MET 1997) Minor failures
The operator mounted a tick tape instead of a thin! First two pass lost.

Calibration (Fredrik T Rantakyro at Tue Dec 9 15:02:47 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name v330c.antab Read the comments in the file to get further information.